Two blocks are connected by a cord passing over a small (frictionless) pulley as shown below. The angle \Theta = 30 degrees, and the mass of the small block is m = 15 kg. If the coefficient of friction between the small block and the inclined plane is 0.25 and the large block is accelerating down at 2 m/s2 , what is the mass, M, of the large block?
Answer #1

mg -T= m(2) -----------1

T - 0.25*9.8*15cos 30 = 30---------------2

adding 1 and 2 we get

mg - 31.82 = 30 + 2m

m = 61.82/(11.8) = 5.23 Kg